IceClaws3309:

The table represents a quadratic function. x y −6 23 −5 8 −4 −1 −3 −4 −2 −1 −1 8 0 23 What is the equation of the function? y = (x + 3)2 − 4 y = (x − 3)2 + 4 y = 3(x + 3)2 − 4 y = 3(x − 3)2 + 4 PLEASE EXPLAIN YOUR ANSWER

Luigi0210:

You could try plugging in coordinates. For example, (0, 23) might easiest since x would just be 0. Using the vertex formula (h, k) , where it's (-3, -4) and plugging in what you can to \[y=a(x-h)^2 +k\] \[ y=a (x+3)^2-4 \] You can narrow it down to either "A" or "C" Just plug in (0, 23) and see which gives you a true statement.

surjithayer:

let\[y=ax^2+bx+c ~...(1)\] when x=0,y=23 23=0+0+c c=23 \[y=ax^2+bx+23~...(2)\] when x=-1,y=8 \[8=a(-1)^2+b(-1)+23\] 8=a-b+23 a-b=8-23=-15 a-b=-15 ...(3) when x=-2,y=-1 \[-1=a(-2)^2+b(-2)+23 \] -1=4a-2b+23 4a-2b=-1-23 4a-2b=-24 divide by 2 2a-b=-12 ...(4) subtract (3)from (4) a=-12+15 a=3 from (3) 3-b=-15 -b=-15-3 b=18 from (1) \[y=3x^2+18x+23\] \[[y=3(x^2+6x+9-9)+23\] \[y=3(x^2+6x+9)-27+23\] y=?